EXTENSIONS OF PERIODIC LINEAR GROUPS WITH FINITE UNIPOTENT RADICAL

Richard E. Phillips† Julianne G. Rainbolt^{1} Jonathan I. Hall^{2} Ulrich Meierfrankenfeld^{3}

1Department of Mathematics, Saint Louis University, Saint Louis, MO 63103

email: rainbolt@slu.edu

2Department of Mathematics, Michigan State University, East Lansing, MI

48824 email: jhall@math.msu.edu

3Department of Mathematics, Michigan State University, East Lansing, MI

48824 email: meier@math.msu.edu

Dedication

This article is published posthumously for the author Richard Phillips. The other authors dedicate this paper to his memory.

Abstract

A group is called p-linear if it is isomorphic to a subgroup of GL(n, K) for some field K of characteristic p and some integer n. Let H be a normal subgroup of G and assume that both H and G/H are periodic and p-linear. In addition, assume that both H and G/H have finite unipotent radicals and that the finite residual of G/H has finite index in G/H. The main result of this article is a proof that under these assumptions G is p-linear. An example is provided showing the result is false if the assumption regarding the finite residual is removed.

1. Results

For finite n, the group G is said to be K-linear of degree n if it is isomorphic to a subgroup of GL(n, K) and is p-linear if it is K-linear for some field K of characteristic p.

The following theorem was proven as [5, Theorem A].

( 1.1) Theorem. Let G be a periodic subgroup of GL(n, K) with trivial unipotent radical. Then for every normal subgroup H of G, the image G/H is Klinear of degree bounded by a function of n.

Here the statement on bounded degree indicates that there is a positive, nondecreasing function f such that the degree of G/H is always less than or equal to f(n).

Wehrfritz [9] gave a second proof of the theorem and was able to weaken the hypotheses by assuming only that H has trivial unipotent radical. A slightly different proof of the theorem is given below (one that in fact includes groups G with finite unipotent radical, as remarked below).

The main concern here is proving a converse result to this theorem. It is not in general the case that an extension of a linear group by a linear group is linear. It is also not clear what shape the best converse to the theorem would take.

Consider the following:

Hypothesis (∗)

G is a locally finite group and M a normal subgroup of G satisfying:

(i) M is a central product A Y L_{1} Y ⋯ Y L_{i} Y ⋯ Y L_{s}, for some s; (ii) each L_{i} is a quasisimple group of Lie type in characteristic p; (iii) A is a periodic abelian p′-group of finite rank r; (iv) G/M is finite.

The rank of a periodic abelian group B is the largest m for which there is a prime q and an elementary abelian q-subgroup C of B having order q^{m}. If p = 0, then s = 0 in (i) and A is periodic abelian of finite rank in (iii).

Each of the groups L_{i} in (i) has, naturally defined, a Lie rank r_{i} and a field of definition K_{i} ≤ \overline{𝔽}_{p}. (See [5, Lemma 9(iv)] or Theorem 2.6 below.) (In characteristic p = 0 we set 𝔽_{p} = ℚ hence \overline{𝔽}_{p} = \overline{ℚ}.) We define the M-rank of G with (∗) to be r +∑_{i=1}^{s} r_{i}. We define the field K^{M} to be the smallest subfield of \overline{𝔽}_{p} that contains each K_{i} and all q^{k}th roots of unity, for every k and all primes q for which A contains a subgroup ℤ_{q^{∞}}. Thus M is K^{M}-linear.

( 1.2) Theorem. If G is a periodic K-linear group of degree n with finite unipotent radical U(G), then G has a characteristic subgroup M for which (G, M) satisfies (∗), where p = Char(K). In this case, M can be chosen so that the M-rank of G is bounded by a function of n and the index |G/M| is bounded by functions of n and |U(G)|.

( 1.3) Theorem. A group G with (∗) is K^{M}-linear of degree bounded by a function of its M-rank and the index |G/M|.

As a corollary, we get Winter’s theorem [10]. The usual proofs are more elementary than this; see [3, Theorem 1.L.2] and [7, Theorem 9.5].

( 1.4) Corollary. If G is a periodic K-linear group with trivial unipotent radical, then G is \overline{𝔽}_{p}-linear, where p is the characteristic of K. In particular, G is countable.

Since uncountable unipotent linear groups of exponent p exist, Winter’s theorem is false without some restriction on the radical. Nevertheless, because of it we can restrict (most) discussion to the question of whether or not a given group is p-linear (that is, linear over some field of characteristic p).

We have a partial converse to Theorem 1.1. For a group G let Res(G) be the finite residual of G, that is the intersection of the subgroups of finite index in G.

( 1.5) Theorem. Let H be a normal subgroup of G and assume that (a) H is a periodic p-linear group with finite unipotent radical;

(b) G/H is a periodic p-linear group with finite unipotent radical;

(c) Res(G/H) has finite index in G/H. Then G is p-linear of degree bounded by a function of deg H, deg G/H and |G/H: Res(G/H)|.

We claim that assumption (c) of the previous theorem is fulfilled if the Hirsch-Plotkin radical of G/H is C̆ernikov. Indeed, let M be as in Theorem 1.2 applied to G/H in place of G. Let E_{∞} be the product of all the infinite L′_{i}s. Since A has finite rank, Res(A) is the divisible part of A. Then (∗) implies that Res(G/H) = Res(A)E_{∞} and so Res(G/H) has finite index in G/H if and only if Res(A) has finite index in A. Note that this fulfilled if the Hirsch Plotkin radical of G/H is C̆ernikov.

( 1.6) Theorem. Let G be a periodic linear group with finite unipotent radical. Let Res(G)^{∗} be the image of Res(G) in Aut(G). Then Out(G) is residually finite and Res(Aut(G)) = Res(G)^{∗}.

The next section contains proofs of Theorems 1.1-1.6. The final section contains two examples. The first example shows that Theorem 1.5 is false under (a) and (b) alone. The other example demonstrates the impossibility in Theorem 1.5 of bounding the representing degree of G in terms of deg G/H and deg H, indeed in terms of the isomorphism class of H and the degree of G/H. The theorem shows that the degree is bounded in terms of the degree of H and the isomorphism class of G/H.

The problem discussed in this article was one that Richard Phillips was working on during the last years of his life. He wrote an initial draft of this article in conjunction with Julianne Rainbolt. After Richard Phillips’ death, Jonathan Hall, Ulrich Meierfrankenfeld, and Julianne Rainbolt completed the revisions of the article. The authors thank Felix Leinen and the referees for helpful remarks on earlier drafts of this article.

We also would like to thank Bert Wehrfritz for noticing that Theorem 1.5 in the published version of this article is false.

2. Proofs

Let G be a locally finite group and π be any set of primes. Then O_{π}(G) denotes the largest normal subgroup of G all of whose elements are π-elements. If G is periodic and linear in characteristic p, then O_{p}(G) is the unipotent radical of G. In the following, E(G) denotes the subgroup of G generated by the components of G, where a component of G is a subnormal quasisimple subgroup.

( 2.1) Theorem. (Schur, [3, Theorem 1.L.1]) Periodic linear groups are locally finite.

The next two lemmas are elementary.

( 2.2) Lemma. If G is K_{0}-linear of degree n and K is a subfield of K_{0} such that |K_{0} : K| is finite, then G is K-linear of degree n|K_{0}: K|.

( 2.3) Lemma. If M is K-linear of degree m and has finite index in G, then G is K-linear of degree m|G: M| via the induced representation.

( 2.4) Lemma. Let H be K-linear and Z a finite normal subgroup of the center Z(H). Then H/Z is K-linear of degree bounded by a function of the degree of H and |Z|.

In particular, if the factors H_{i} are K-linear, then a central product H_{1} Y H_{2} = (H_{1} × H_{2})/Z over a finite central subgroup Z is K-linear of degree bounded by a function of the degrees of the H_{i} and |Z|.

Proof. As the direct product of two K-linear groups is K-linear with degree equal to the sum of the two degrees, this an immediate consequence of [5, Proposition 3(ii)].

( 2.5) Lemma. Let B be a periodic abelian group and let A be the divisible hull of B. Suppose K contains an n-root of unity whenever A contains an element of order n. Let V be faithful finite dimensional KB-module. If the unipotent radical of B is trivial, then V can be extended to a faithful KA-module with End_{KA}(V) = End_{KB}(V).

Proof. As V is finite dimensioal and B has trivial unipotent radical, B has finite rank. Note that every non-trivial subgroup of A intersects B and so any extension of V to a KA is faithful. Any direct summand of B is contained in a direct summand of A and so by induction on the rank of B we may assume that B = ⟨b⟩ is cylic and hence A is locally cyclic. Let λ be an eigenvalue for b on V.

Then there exists an homomorphism ϕ_{λ} : A → (K^{♯}, ·) with ϕ(b) = λ. From the assumption, V is the direct sum of the eigenspaces for b. Define v^{a} = ϕ_{λ}(a)v whenever a ∈ A and v is in the eigenspace corresponding to λ.

A proof of the following theorem can be found at [4, Lemmas 15.6, 15.10, and Theorem 15.12]. It is very similar to [5, Proposition A] and [8, 1.2].

( 2.6) Theorem. Let G be a periodic linear group of degree n over a field in characteristic p and having trivial unipotent radical.

Let { L_{i} | 1 ≤ i ≤ t } consist of all components of G that have Lie type in characteristic p. The central product E_{L,p}(G) = L_{1} Y ⋯ Y L_{t} is characteristic in G, and t ≤ n/2. Furthermore, G has a characteristic abelian subgroup A such that the subgroup M(G) = A Y E_{L,p}(G) is characteristic in G and has finite index bounded by a function of n.

Proof of Theorem 1.2.

G is locally finite by Theorem 2.1.

As the unipotent radical U = U(G) is finite, C_{G}(U) is a characterisitic subgroup of G of finite index bounded by a function of |U|. Therefore we may assume that U ≤ Z(G).

By Lemma 2.4, Ḡ = G/U is K-linear with trivial unipotent radical of degree bounded by a function of |U| and n.

Let N̄ = M(Ḡ) be the subgroup Ā Y L̄_{1} Y ⋯ Y L̄_{s} of Theorem 2.6. Let L_{i} be the derived group of the preimage of L̄_{i} in G, A the p′-part of the preimage of Ā, and M = A Y L_{1} Y ⋯ Y L_{s}. Then M is characteristic in G, and the pair (G, M) satisfies (∗).

The various rank(L_{i}) are bounded by a function of n [5, Lemma 9(b)(i)]. Since A is p′ and periodic, it has finite rank bounded by a function of n by Maschke’s Theorem.

Proof of Theorem 1.3.

By (iv) and Lemma 2.3, we may assume G = M. Each L_{i} is K^{M}-linear by assumption and has a center that is finite of order bounded by a function of the Lie rank of L_{i} ([5, Lemma 10(viii)] or Lemma 2.8). Thus E(G) = L_{1} Y ⋯ Y L_{s} is K^{M}-linear by Lemma 2.4 and has finite center of order bounded in terms of the Lie ranks of the L_{i}. By definition, (K^{M})^{×} contains a copy of ℤ_{q^{∞}} whenever A has infinite q-part, so A is K^{M}-linear of degree equal to its rank r. A second application of Lemma 2.4 then proves that G = M = A Y E(G) is K^{M}-linear, as desired. Furthermore, its degree is controlled as described.

Proof of Corollary 1.4.

As G is periodic and K-linear, it has (∗) by Theorem 1.2 and so is \overline{𝔽}_{p}-linear by Theorem 1.3.

Proof of Theorem 1.1.

By Theorem 1.2, G has (∗) for a normal subgroup M, and the M-rank of G and |G/M| are both bounded by functions of n. In Ḡ = G/H, set M̄ = MH/H. Then the pair (Ḡ, M̄) inherits (∗), so by Theorem 1.3 Ḡ is linear of degree bounded in terms of the M̄-rank of Ḡ (at most the M-rank of G) and |Ḡ: M̄| (≤ |G: M|). Thus the degree of G/H is bounded by a function of n, as desired.

Remark. The same proof actually gives something slightly stronger than Theorem 1.1. We need only require that the unipotent radical of G be finite, in which case the representation degree of G/H is bounded by a function of n and of the order of the radical.

( 2.7) Lemma. Let B be a class 2 nilpotent group with B′ ≤ H ≤ Z(B). Assume that B/H is divisible and periodic. Then B is abelian.

Proof. Let b ∈ B and choose an integer n with b^{n} ∈ H. Then B/C_{B}(b) ≅ [B, b] has exponent dividing n. But H ≤ C_{B}(b) and so B/C_{B}(b) is divisible. Thus B = C_{B}(b) and B is abelian.

( 2.8) Lemma. Let periodic N have a central subgroup H such that N̄ = N/H is a central product M̄_{1} Y ⋯ Y M̄_{t} of finitely many infinite quasisimple groups M̄_{i} of Lie type in characteristic p. Then N = H Y E(N), where E(N) = N_{1} Y ⋯ Y N_{t} has finite center of deg N̄-bounded order, and is a central product of quasisimple groups N_{i} of Lie type with N̄_{i} = M̄_{i}. Moreover, E(N) is p-linear of deg N̄-bounded degree.

Proof. By the Three Subgroups Lemma, Z_{2}(N) = Z(N). Also N′ = [N′H, N′H] ≤ N′′; so N′ is perfect, and N = H Y N′. Indeed N/H = N′H/H ≃ N′/N′ ∩ H has image N′/Z(N′). Therefore N′ = E(N), the central product of the components N_{i} = (M_{i})′, the derived subgroups of the preimages M_{i} of the various M̄_{i}.

Each simple periodic infinite Lie type group L is a direct limit of finite simple groups of the same Lie type, and so any element of its multiplier occurs already within the multiplier of some finite subgroup of Lie type. Exceptional multipliers for finite Lie type groups occur only over small fields, and the canonicial multipliers come from the natural or spin representations of fixed degree bounded in terms of the Lie rank [2]. Thus the multiplier of L is finite of order bounded by the rank of L and comes from a representation of degree also bounded by the rank. (See also [5, Lemma 10].)

Proof of Theorem 1.6. By 1.2 there exists a characteristic subgroup M of G fulfilling (∗). Let E be the product of the infinite L_{i}. Using Lemma 2.8 we may choose A to be a characteristic subgroup of M. Then EA is a characterisic subgroup of finite index. Let H = Res(G) and note that H = Res(A)E and Res(A) is the largest divisible subgroup of A. For g ∈ G let g^{∗} ∈ Aut(G), h → h^{g}.

We first treat the case where E = 1. Let π be the set of prime divisors of |G/A| and let D be the largest π-divisible subgroups of A. Note that A has finite rank, O_{π′}(A) ≤ D and H = Res(A) ≤ D. Therefore A/D and G/D are finite π-groups. Let n = |G/D| and for a positive integer i let D_{i} = { d ∈ D | d^{i} = 1 }. Then D_{i} is finite. The following cohomological argument is taken from [1, Propostions 3.7.5]. Since D is n divisible, the following is a short exact sequence:

1 ⟶ D_{n} ⟶ D ⟶^{a→a^{n}} D ⟶ 1

Hence we also obtain the long exact sequence

… ⟶ H^{m}(G/D, D_{n}) ⟶ H^{m}(G/D, D) ⟶^{α→nα} H^{m}(G/D, D) ⟶ …

Suppose m > 0. By [1, Proposition 3.6.17] the map α → nα factors through the restriction map to the trivial subgroup of G/D and so is 0. Thus

H^{m}(G/D, D_{n}) → H^{m}(G/D, D)

is onto. That is every cocyle for G/D in D arises from a cocycle of G/D in D_{n}.

For m = 2 we conclude that there exists T ≤ G with G = DT and D ∩ T = D_{n}. For m = 1 and G replaced by G/D_{n} we see that for all T̃ ≤ G with G = DT̃ and G ∩ T̃ = D_{n}, there exists d ∈ D with T̃ ≤ T^{d}D_{n^{2}}. Put R = T D_{n^{2}}. Let α ∈ Aut(G) and choose d ∈ D with T^{α} ≤ T^{d}D_{n^{2}}. Then R^{α} = R^{d} and so Aut(G) = N_{Aut(G)}(R)D^{∗}.

Put F = C_{Aut(G)}(R). Since R is finite we get that F D^{∗} has finite index in Aut(G). Note that C_{F} (D) ≤ C_{F} (G) = 1 and so F is isomorphic to a subgroup of Aut(D). Since D is generated by its finite characteristic subgroup, Aut(D) is residually finite. Thus F is residually finite. Note that F ∩ D^{∗} ≤ C_{F} (D) = 1. Hence F D^{∗}/D^{∗} is residually finite and since this group has finite index in Aut(G)/D^{∗}, Aut(G)/D^{∗} is residually finite. Moreover, G^{∗}/D^{∗} is finite and so Aut(G)/G^{∗} is residually finite.

Let p be a prime. Then O_{p′}(D)H has finite index in D and so F O_{p′}(D)^{∗}H^{∗} has finite index in Aut(G). Also F O_{p′}(D)^{∗}H^{∗}/O_{p′}(D)^{∗}H^{∗} ≅ F is residually finite and so Res(Aut(G)) ≤ O_{p′}(D)^{∗}H^{∗}. Since this holds for all primes, Res(Aut(G)) ≤ H^{∗}. But H^{∗} has no proper subgroups of finite index and so Res(Aut(G)) = H^{∗}.

Hence 1.6 holds if E = 1. In the general case, we have E = L_{1}L_{2} … L_{k} where L_{i} is a group of Lie type ^{σ_{i}}Φ_{i} over an infinite field K_{i}. Let F_{i} be a finite subfield of K_{i} such that if σ_{i} ≠ 1, σ_{i} acts nontrivially on F_{i}. Let L_{i}(F_{i}) be the group of Lie type ^{σ_{i}}Φ_{i} over the field F_{i} naturally embedded into L_{i}. We choose the F_{i} such that the Schur multipliers of L_{i} and L_{i}(F_{i}) are identical. Let X = ∏_{i=1}^{k} L_{i}(F_{i}) ≤ E. Since G/C_{G}(E)E is finite we we can choose the F_{i}’s such that C_{G}(X) ≤ C_{G}(E). Note that L_{i}(F_{i}) is normalized by any field or graph automorphism of L_{i}. Moreover, our condition on the Schur multiplier ensures that every diagonal automorphism of L_{i} can be written as a product of an inner automorphism and an automorphism normalizing L_{i}(F_{i}).

Put B = N_{G}(X) and F = N_{Aut(G)}(X). By the preceeding discussion G = EB and Aut(G) = F E^{∗}. Since B has no infinite component, we can apply the E = 1-case to B. We conclude that both F/Res(B)^{∗}C_{F} (B) and F/B^{∗}C_{F} (B) are residually finite. Note that

Aut(G)/G^{∗} = F G^{∗}/G^{∗} ≅ F/F ∩ G^{∗} = F/B^{∗}

and so Res(Aut(G)/G^{∗}) ≤ C_{F} (B)G^{∗}/G^{∗}. Also C_{G^{∗}}(B) ≤ C_{G^{∗}}(X) ≤ C_{G^{∗}}(E). Since BE = G we conclude that C_{G^{∗}}(B) = 1. In particular, G^{∗} ∩ C_{F} (B) = 1. Thus

⋂ C_{F} (B)G^{∗} = (⋂ C_{F} (B))G^{∗} = G^{∗}

where the intersection is taken over all the eligible F_{i}. Thus Res(Aut(G)/G^{∗}) = 1.

Since B/C_{B}(X), C_{G}(X)/C_{G}(E) and C_{G}(E)/A all are finite we conclude that Res(B) = Res(A) ≤ F and H = Res(B)E. So C_{F} (B)Res(B)^{∗} ≤ F,

F ∩ (C_{F} (B)Res(B)^{∗}E^{∗}) = C_{F} (B)Res(B)^{∗}(F ∩ E^{∗}) = C_{F} (B)Res(B)^{∗}(B ∩ E)^{∗}

and

Aut(G)/C_{F} (B)H^{∗} ≅ F E^{∗}/C_{F} (B)Res(B)^{∗}E^{∗} ≅ F/C_{F} (B)Res(B)^{∗}(B ∩ E)^{∗}.

Since B∩E is finite and F/C_{F} (B)Res(B)^{∗} is residually finite, Res(Aut(G)) ≤ C_{F} (B)H^{∗}. Intersecting over the various choices for F_{i} gives Res(Aut(G)) ≤ H^{∗}. Also H has no proper subgroup of finite index and so Res(Aut(G)) = H^{∗}.

Proof of Theorem 1.5.

By 2.3 we may assume that G/H = Res(G/H) and so G/H has no nontrivially residually finite quotient. By Theorem 1.6 G/C_{G}(H)H is residually finite and so G = C_{G}(H)H. Thus C_{G}(H)/Z(H) ≅ G/H. Since Res(G)/H = G/H we conclude that C_{G}(H) = A/Z(H) × L/Z(H) where A/Z(H) is divisible abelian and L/Z(H) is the central product of infinite groups of Lie type. By 2.8, L = E(L)Z(H), E(L) is p-linear of deg G/H-bounded degree and |Z(L′)| is deg G/H-bounded. So by 2.4 there exists a faithful \overline{𝔽}_{p}E(L)H module V of dimension bounded by a function of deg H and deg G/H.

By 2.7 A is abelian. Let A_{0} ≤ Res(A) with Res(A) = Res(Z(H)) × A_{0}. Let A_{1} be a divisible hull for Z(H) ∩ A_{0} in A_{0}. Then there exists A_{2} ≤ A_{0} with A_{0} = A_{2} × A_{1}. Moreover, G = A_{2} × (A_{1}E(L)H. According to 2.5 we can extend the \overline{𝔽}_{p}(Z(H) ∩ A_{0})-module W to a module for A_{1}. So W becomes an \overline{𝔽}_{p}G module with A_{2} acting trivially. Take the direct sum of W with a faithful G/H module we obtain a faithful \overline{𝔽}_{p}G-module of dimension bounded in terms of deg H and deg G/H.

3. Examples

( 3.1) Lemma. There is an infinite sequence of primes

q_{1} < p_{1} < ⋯ < q_{i} < p_{i} < ⋯

for i ∈ ℤ^{+}, with q_{i}| p_{i} − 1.

Proof. The proof is by induction on i. Let q_{1} be any prime number. A theorem of Dirichlet [6, p. 250] asserts that the sequence {1 + kq_{1} | k = 1, 2, …} contains an infinite number of primes. Let p_{1} be any such prime. Then q_{1}|(p_{1} − 1). Now let q_{2} be a prime greater than p_{1}.

Suppose that we have chosen q_{1}, …, q_{s+1} and p_{1}, …, p_{s} such that

q_{1} < p_{1} < ⋯ < q_{s} < p_{s} < q_{s+1}

Now choose the prime p_{s+1} in the sequence {1 + kq_{s+1} | k = 1, 2, …} and the prime q_{s+2} with q_{s+2} > p_{s+1}.

( 3.2) Example.

We first show that Theorem 1.5 is false when only (a) and (b) are assumed (even when (a) and (b) are strengthened to require that the unipotent radicals be trivial).

For p_{i}, q_{i} (i ∈ ℤ^{+}) as in Lemma 3.1, let S_{i} be cyclic of order p_{i} and R_{i} cyclic of order q_{i}. Let F_{i} = S_{i}.R_{i} be the Frobenius group of order p_{i}q_{i}. Then F^{k} = ⊕ ∑_{i=1}^{k} F_{i} has representation degree at least 2k in each characteristic, even though it is the extension S^{k}.R^{k} of cyclic S^{k} = ⊕ ∑_{i=1}^{k} S_{i} of degree 1 over any prime not in the sequence (and degree at most 2 in general) by cyclic R^{k} = ⊕ ∑_{i=1}^{k} R_{i} also of degree 1 over any prime not in the sequence (and degree at most 2 in general).

Thus F^{∞} = ⊕ ∑_{i} F_{i} = lim_{k} F^{k} is not linear in any characteristic even though it is the extension of rank 1 linear S^{∞} = ⊕ ∑_{i} S_{i} = lim_{k} S^{k} by rank 1 linear R^{∞} = ⊕ ∑_{i} R_{i} = lim_{k} R^{k}. The residual core of F^{∞}/S^{∞} is trivial and has infinite index in F^{∞}/S^{∞}.

( 3.3) Example.

The groups F^{k} of the first example show that the degree of linear G can not be bounded in terms of the degree, deg H, of the normal subgroup H and deg G/H. The next example will show that the degree of linear G can not in general be bounded in terms of deg G/H and the isomorphism type of H.

Fix a characteristic p and a prime q ≠ p. Let, for n ∈ ℤ^{+},

E(n) = ⟨x, y, z | x^{q^{n}} = y^{q^{n}} = z^{q^{n}} = 1, [x, y] = z, [x, z] = [y, z] = 1 ⟩ .

Let E^{∗}(n) be the central product of E(n) and ℤ_{q^{∞}} with z identified with an element of order q^{n} of ℤ_{q^{∞}}. Then H = ℤ_{q^{∞}} is p-linear of degree 1 and E^{∗}(n)/H ≃ Z_{q^{n}} × Z_{q^{n}} is p-linear of degree 2.

We claim that the degree of E(n) (and hence of E^{∗}(n)) in characteristic p is q^{n}. Abelian W = ⟨x, z⟩ has index q^{n}, so this is an upper bound on the degree. Consider the restriction of a faithful representation to W. There is some degree 1 constituent ζ with W = ⟨w, z⟩ = C_{E(n)}(w), where ⟨w⟩ = ker(ζ). Then by Clifford’s Theorem the ζ^{y^{i}}, for i = 0, …, q^{n} − 1, are distinct constituents of the restriction.

References

[1] D.J. Benson, Representation and Cohomology I, Cambridge University Press,Cambridge, 1991.

[2] R.L. Griess, Jr., Schur multipliers of the known finite simple groups II, in: Finite Groups, Proc. Symp. Pure Math. 37 (1980), 279–282.

[3] O.H. Kegel and B.A.F. Wehrfritz, Locally Finite Groups, North Holland, Amsterdam, 1973.

[4] U. Meierfrankenfeld, Lecture notes on locally finite groups, 1997 (available from http://www.math.msu.edu/∼meier/)

[5] R.E. Phillips and J.G. Rainbolt, Images of periodic linear groups, Arch. Math. 71 (1998), 97–106.

[6] P. Ribenboim, The Book of Prime Number Records, Springer-Verlag, Berlin-Heidelberg-New York, 1988.

[7] B.A.F. Wehrfritz, Infinite Linear Groups, Springer-Verlag, Berlin-Heidelberg-New York, 1973.

[8] B.A.F. Wehrfritz, Locally finite normal subgroups of absolutely irreducible skew linear groups, J. London Math. Soc. (2) 32 (1985), 88–102.

[9] B.A.F. Wehrfritz, Periodic normal subgroups of linear groups, Arch. Math. 71 (1998), 169–172.

[10] D.J. Winter, Representations of locally finite groups, Bull. Amer. Math. Soc. 74 (1968), 145–148.